If the gravitational force is 12663 N between a satelite with a mass of 11 x 10^2 kg and a planet with a mass of 2 x 10^20 kg, find the distance between them
Answer #1

here,

mass of satellite , m1= 110 kg

mass of planet , m2 = 2 * 10^20 kg

let the distance between them be r

the gravitational force , F = G * m1 * m2 /r^2

12663 = 6.67 * 10^-11 * 110 * 2 * 10^20 /r^2

solving for r

r = 1.08 * 10^4 m

the distance between them is 1.08 * 10^4 m